OpenStudy (anonymous):

what is the solution to -19*^7sqrt15x+8=-38 I think its 8 not sure though

OpenStudy (anonymous):

Yes, you are correct. Here is what I got. \[-19\sqrt[7]{15x+8}=-38\] \[\sqrt[7]{15x+8}=\frac{-38}{-19}=2\] \[\left( \sqrt[7]{15x+8} \right)^7=2^7\] \[15x+8=128\] \[15x=120\] \[x=8\]

OpenStudy (anonymous):

Believe in yourself. :)